# BEVDiffuser: Plug-and-Play Diffusion Model for BEV Denoising with Ground-Truth Guidance

### v1

@@ 1 Introduction
Despite recent advancements in BEV generation, the issue of noise in these BEV representations remains largely unresolved. Generated BEV representations are inherently noisy due to the imperfections of acquisition sensors such as camera and LiDAR, as well as the limitations in the learning process [41, 12]. The noise from acquisition sensors introduces inaccuracies, including imprecise localization of object boundaries in BEV feature maps, which degrades performance in downstream tasks. Additionally, in the absence of direct supervision, BEV representations are typically optimized only for downstream task performance, leading to potential biases within the BEV feature maps. Generative models, particularly diffusion models, are well-suited to address this challenge due to their powerful denoising capabilities [26, 27, 23]. Diffusion models have demonstrated remarkable success in image and video generation [22, 23, 1], and recent studies have extended their applicability to tasks such as image classification and object detection [13, 3]. Leveraging diffusion models to denoise and enhance BEV representations holds significant potential for improving the robustness and accuracy of BEV-based downstream tasks.
@@
BEVDiffuser, as a flexible plug-and-play module, can be seamlessly incorporated into any BEV model. In this study, we conduct an extensive evaluation of BEVDiffuser on four widely adopted state-of-the-art BEV models using the challenging nuScenes [2] dataset. The experimental results demonstrate BEVDiffuser’s exceptional denoising capabilities, which enable significant enhancements to existing BEV models, demonstrated by notable improvements of $12.3\%$ in mAP and $10.1\%$ in NDS for 3D object detection. Additionally, our experiments show that BEVDiffuser substantially improves performance in long-tail object detection and under challenging weather and lighting conditions, highlighting its ability to produce more accurate and robust BEV representations. Furthermore, BEVDiffuser also shows high-quality BEV generation capabilities from pure noise with layout conditioning, which can pave the way for large-scale data collection to advance autonomous driving. Qualitative visualizations further validate the observed quantitative improvements.
@@
  • We conduct extensive experiments on the challenging nuScenes dataset. The results show that BEVDiffuser has strong denoising capabilities, and significantly improves BEV models both quantitatively and qualitatively.
  • BEVDiffuser demonstrates enhanced robustness in long-tail cases and under adversarial weather and lighting conditions, while also showcasing powerful BEV generation capabilities.
@@ 2.1 BEV Feature Map
BEV feature generation, which fuses multi-view sensor inputs into a top-down feature map, provides comprehensive and holistic information about the scene. It has recently gathered immense interest, as it plays a key role in the rapid advancement of autonomous driving [7, 10]. Many works focus on learning BEV features from cameras alone, which can be broadly categorized into two main approaches: geometry-based methods, represented by Lift-Splat-Shoot (LSS) [21], and transformer-based methods, exemplified by BEVFormer [15]. LSS [21] generates BEV feature maps from multi-view images by leveraging the estimated depth distribution. BEVDepth [14] further improves LSS by explicitly supervising the depth estimations. The BEV feature maps generated by LSS are used by BEVDet [9] and BEVDet4D [8] for 3D object detection. In contrast, transformer-based methods leverage powerful attention mechanism to extract attended image features for BEV generation. BEVFormer has gained significant attention as it captures both spatial and temporal information through spatial cross-attention and temporal self-attention mechanisms, respectively [15]. Subsequently, numerous follow-up works such as the adaptation of modern image backbones and the inclusion of additional supervisions have been proposed to further enhance the quality of generated feature maps [32]. Another line of work presents strategies to fuse multi-modal sensor inputs for more robust BEV feature generation [17, 19, 18]. BEVFusion [19] is a representative work that introduces a unified framework for camera and LiDAR sensors by combining multi-modal features in BEV space. In contrast to these works, we propose a plug-and-play diffusion model designed to enhance the BEV feature maps by denoising the intrinsic noise from both the acquisition sensors and the learning process.
@@ 2.2 Diffusion Model Enhanced BEV
Diffusion models [26, 28, 6] are a class of generative models that have demonstrated impressive performance and stability compared to other generative approaches, such as GANs and VAEs. While diffusion models have been primarily used for generative tasks, such as image generation [23, 36, 22, 38] and video generation [25, 1, 31, 29], their applications to downstream tasks such as image classification [13], object detection [3], semantic segmentation [16], motion prediction [11], and 3D shape generation [39] have recently been investigated.
Only a few approaches have been proposed to use diffusion models for enhancing the BEV feature maps [12, 41], which is the focus of this study. Specifically, DiffBEV [41] applies a conditional diffusion model to progressively refine the noisy BEV feature maps, using the learned features as conditions. The denoised BEV is then fused with the original BEV to perform downstream tasks. Similarly, DifFUSER [12] leverages a diffusion model for better sensor fusion. It enhances the fused features obtained from camera and LiDAR sensors by denoising them conditioned on partial camera and LiDAR features during run time. Both approaches demonstrate the potential of diffusion models for denoising and enhancing BEV feature maps. However, unlike our BEVDiffuser, these approaches rely on noisy information as conditions to guide the denoising process which is less effective. Moreover, they require multiple passes through their integrated diffusion model during inference, making them computationally expensive for latency-critical real-world applications like autonomous driving..
@@ 3.1 Preliminary
Diffusion Model. Diffusion model has been proposed for data generation. To generate a data from a random noise $\mathcal{N}(\bm{0},\bm{I})$, diffusion model first destroys structure in a data distribution by gradually adding noise to the data samples, then it learns a reverse denoising process to restore the data structure. Specifically, given a timestep $t\sim Uniform(\{1,...,T\})$, it adds $t$-step noise to the data sample $\bm{x}_{0}$ and get noisy sample $\bm{x}_{t}=\sqrt{\bar{\alpha_{t}}}\bm{x_{0}}+\sqrt{1-\bar{\alpha_{t}}}\bm{\epsilon}_{t}$, where $\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\bm{I})$ and $\bar{\alpha_{t}}=\prod_{i=1}^{t}\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{t})$ in which $\beta_{t}\in(0,1)$ is a hyperparameter that controls the noise strength. Diffusion model then learns a function $f_{\theta}(\bm{x}_{t},t)$, typically modeled by a U-Net [24], to estimate the $\bm{\epsilon}_{t}$ by minimizing the diffusion loss $\mathcal{L}_{diffusion}=\mathbb{E}_{\bm{x}_{0},\bm{\epsilon}_{t},t}\parallel\bm{\epsilon}_{t}-f_{\theta}(\bm{x}_{t},t)\parallel_{2}^{2}$. Once the function $f_{\theta}(\bm{x}_{t},t)$ is well learned, a new data $\bm{x}_{0}$ can be generated from the random noise $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$ through the iterative sampling process $\bm{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha_{t}}}}\bm{\epsilon}_{t})+\sigma_{t}\bm{z}$, where $\bm{\epsilon}_{t}$ is estimated by the learned function $f_{\theta}(\bm{x}_{t},t)$, $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$, and $\sigma_{t}$ is usually set to $\beta_{t}$ or scaled form of $\beta_{t}$.
More recently, to have a control over the denoising process and generate data of interest, conditional diffusion model with classifier-free guidance is often used because of its efficiency [5]. In particular, a condition $y$ is fed into $f_{\theta}$ with a certain probability during training to get the conditional estimation of the noise $\bm{\epsilon}_{t}$. During sampling, the noise $\bm{\epsilon}$ is estimated by $(1+w)f_{\theta}(\bm{x}_{t},t,y)-wf_{\theta}(\bm{x}_{t},t,y=\phi)$ with the weight $w$ being set to balance the conditional and unconditional estimations.
@@ 3.2 BEVDiffuser with Ground-Truth Guidance
Training. In the absence of ground-truth BEV feature map $\bm{x}_{0}$, we add noise $\bm{\hat{\epsilon}}_{t}$ to the predicted BEV $\bm{x}_{t_{0}}(0\leq t_{0}\ll T)$ to get $\bm{x}_{t}$. In this case, we don’t have the access to the true noise $\bm{\epsilon}_{t}$, which is supposed to be added to $\bm{x}_{0}$ to generate $\bm{x}_{t}$. As a result, instead of using $f_{\theta}$ to estimate the unknown $\bm{\epsilon}_{t}$, we propose to optimize $f_{\theta}$ towards $\bm{x}_{0}$. Since $\bm{x}_{t_{0}}$ is already a good estimation of $\bm{x}_{0}$ with bounded task errors, we first optimize $f_{\theta}$ towards $\bm{x}_{t_{0}}$ by minimizing the diffusion loss $\mathcal{L}_{diffusion}$ defined in Equation 1. To further improve the estimation accuracy, we attach task heads to consume the outputs of $f_{\theta}$ and generate task-specific predictions. In this way, $f_{\theta}$ can also be optimized through the task-specific loss $\mathcal{L}_{task}$. To sum up, we adopt the weighted sum of both losses as the overall loss $\mathcal{L}_{total}^{\scriptscriptstyle diff}$ to train $f_{\theta}$. Equation 2 defines the loss $\mathcal{L}_{total}^{\scriptscriptstyle diff}$ where $\lambda$ denotes a weight.
@@
Sampling. We adopt classifier-free guidance in sampling process where we interpolate between conditional and unconditional outputs of $f_{\theta}$ to get the final estimation of $\bm{x}_{0}$ as Equation 3 calculates. The unconditional estimation of $\bm{x}_{0}$ is obtained by replacing the conditioning layout $l$ with the empty layout $l_{\phi}=\{o_{0},o_{p},...,o_{p}\}$ that only contains points $o_{p}$. We then derive $\bm{\epsilon}_{t}$ from the estimated $\bm{x}_{0}$ by Equation 4 for the iterative sampling process.
@@ 4 Experiments
We validate BEVDiffuser on 3D object detection task, the most common task used to evaluate the effectiveness of the learned BEV feature maps [15, 19, 12, 41]. 3D object detection is critical in autonomous driving that requires both semantic and geometric understanding of the environment to identify and locate objects in a 3D space. In this section, we first introduce our experimental setting in Sec. 4.1. In Sec. 4.2, we showcase the capacity of BEVDiffuser in denoising and generating BEV feature maps. We further demonstrate plug-and-play performance of BEVDiffuser in Sec. 4.3 by comparing BEVDiffuser enhanced BEV models with their baseline counterparts.
@@ 4.3 Plug-and-Play Performance of BEVDiffuser
It is worth highlighting that BEVDiffuser brings performance enhancement to BEV models at no cost of any additional adaptation efforts or computational overhead. As a training-only plug-in, BEVDiffuser is removed at deployment, leaving an enhanced BEV model with the architecture unchanged, which is then used for testing. As a result, our BEVDiffuser enhanced models share the same network size and latency as their baseline counterparts which are summarized in Tab. 2. Unlike previous work [12, 41] that need to pass their integrated diffusion models multiple times to denoise the BEV feature maps on-the-fly [^1] We couldn’t compare latency with them as they are not open-sourced. , our method is more flexible and superior in latency-critical applications like autonomous driving.
@@ 5 Conclusion and Future Work
In this work, we present BEVDiffuser, a novel diffusion model that denoises BEV feature maps using ground-truth guidance. BEVDiffuser consists of a U-Net model trained on BEV feature maps generated by existing BEV models. The U-Net model predicts clean BEV feature maps conditioned on the ground-truth object layout, which then derives the denoising process. BEVDiffuser can be used as a training-only plug-and-play module to enhance the existing BEV models by providing denoised BEV feature maps as additional supervision to BEV predictions. Extensive experiments on challenging nuScenes dataset demonstrate BEVDiffuser’s exceptional denoising and generation capabilities, resulting in significant improvements to existing BEV models, without the need for architectural changes or additional computational overhead. Moreover, results on long-tail object detection and under challenging weather and lighting conditions further confirm the efficacy of BEVDiffuser in improving the BEV quality.
In future work, we plan to investigate potential applications of BEVDiffuser for other autonomous driving tasks, such as motion prediction and data augmentation for corner case scenarios. Moreover, while our study presents a novel method for BEV, it also opens pathways for leveraging diffusion models in broader 3D perception research areas, such as sparse query-based methods for autonomous driving.

### v2

@@ 1 Introduction
(a) BEVFormer [15]
(b) BEVDiffuser
(c) LiDAR Top View
Figure 1: Comparisons of BEV feature maps: (a) generated by BEVFormer (tiny) [15], (b) denoised by BEVDiffuser in $5$ steps. Channel-wise features are averaged for visualization. BEVDiffuser denoises and substantially enhances the BEV feature maps.
@@
Despite recent advancements in BEV generation, the issue of noise in these BEV representations remains largely unresolved. Generated BEV representations are inherently noisy (see Fig. 1(a)) due to the imperfections of acquisition sensors such as camera and LiDAR, as well as the limitations in the learning process [41, 12]. The noise from acquisition sensors introduces inaccuracies, including imprecise localization of object boundaries in BEV feature maps, which degrades performance in downstream tasks. Additionally, in the absence of direct supervision, BEV representations are typically optimized only for downstream task performance, leading to potential biases within the BEV feature maps. Generative models, particularly diffusion models, are well-suited to address this challenge due to their powerful denoising capabilities [27, 28, 24]. Diffusion models have demonstrated remarkable success in image and video generation [23, 24, 1], and recent studies have extended their applicability to tasks such as image classification and object detection [13, 3, 20]. Leveraging diffusion models to denoise and enhance BEV representations holds significant potential for improving the robustness and accuracy of BEV-based downstream tasks.
@@
BEVDiffuser, as a flexible plug-and-play module, can be seamlessly incorporated into any BEV model. In this study, we conduct an extensive evaluation of BEVDiffuser on four widely adopted state-of-the-art BEV models using the challenging nuScenes [2] dataset. The experimental results demonstrate BEVDiffuser’s exceptional denoising capabilities (see Fig. 1(b)), which enable significant enhancements to existing BEV models, demonstrated by notable improvements of $12.3\%$ in mAP and $10.1\%$ in NDS for 3D object detection. Additionally, our experiments show that BEVDiffuser substantially improves performance in long-tail object detection and under challenging weather and lighting conditions, highlighting its ability to produce more accurate and robust BEV representations. Furthermore, BEVDiffuser also shows high-quality BEV generation capabilities from pure noise with layout conditioning, which can pave the way for large-scale data collection to advance autonomous driving. Qualitative visualizations further validate the observed quantitative improvements.
@@
  • Extensive experiments on the nuScenes dataset demonstrate that BEVDiffuser possesses strong BEV denoising and generation capabilities, significantly enhances BEV models both quantitatively and qualitatively, and exhibits improved robustness in long-tail cases and adverse weather and lighting conditions.
@@ 2.1 BEV Feature Map
Camera-only BEV feature generation works can be broadly categorized into two main approaches: geometry-based methods, represented by Lift-Splat-Shoot (LSS) [22], and transformer-based methods, exemplified by BEVFormer [15]. LSS [22] generates BEV feature maps from multi-view images by leveraging the estimated depth distribution, followed by [14, 9, 8]. In contrast, transformer-based methods utilize powerful attention mechanism to extract attended image features for BEV generation. BEVFormer [15] and its follow-up BEVFormerV2 [33] have gained significant interest as they capture both spatial and temporal information through spatial cross-attention and temporal self-attention mechanisms, respectively. Another line of work presents strategies to fuse multi-modal sensor inputs for more robust BEV feature generation [17, 19, 18]. BEVFusion [19] is a representative work that introduces a unified framework for camera and LiDAR sensors by combining multi-modal features in BEV space. In contrast to these works, we propose a plug-and-play diffusion model designed to enhance the BEV feature maps by denoising the intrinsic noise from both the acquisition sensors and the learning process.
@@ 2.2 Diffusion Model Enhanced BEV
Diffusion models are a class of generative models that have demonstrated impressive performance and stability [27, 29, 6]. While diffusion models have been primarily used for generative tasks, such as image generation [24, 37, 23, 39] and video generation [26, 1, 32, 30], their applications to downstream tasks such as image classification [13], object detection [3], semantic segmentation [16], and motion prediction [11] have recently been investigated.
Only a few approaches have been proposed to use diffusion models for enhancing the BEV feature maps [12, 41], which is the focus of this study. Specifically, DiffBEV [41] applies a conditional diffusion model to progressively refine the noisy BEV feature maps, using the learned features as conditions. The denoised BEV is then fused with the original BEV to perform downstream tasks. Similarly, DifFUSER [12] leverages a diffusion model for better sensor fusion. It enhances the fused features obtained from camera and LiDAR sensors by denoising them conditioned on partial camera and LiDAR features during run time. Both approaches demonstrate the potential of diffusion models for denoising and enhancing BEV feature maps. However, unlike our BEVDiffuser, these approaches rely on noisy information as conditions to guide the denoising process which is less effective. Moreover, they require multiple passes through their integrated diffusion model during inference, making them computationally expensive for latency-critical real-world applications like autonomous driving.
@@ 3.1 Preliminary
Diffusion Model. Diffusion model learns to generate data from random noise $\mathcal{N}(\bm{0},\bm{I})$ by first destroying the structure of a data distribution through gradual addition of noise to the data samples, and then learning a reverse denoising process to restore the data structure. Specifically, given a timestep $t\sim Uniform(\{1,...,T\})$, it adds $t$-step noise to a data sample $\bm{x}_{0}$ to get a noisy sample $\bm{x}_{t}$:
$\bm{x}_{t}=\sqrt{\bar{\alpha_{t}}}\bm{x_{0}}+\sqrt{1-\bar{\alpha_{t}}}\bm{\epsilon}_{t},$ (1)
where $\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\bm{I})$ and $\bar{\alpha_{t}}=\prod_{i=1}^{t}(1-\beta_{t})$ in which $\beta_{t}\in(0,1)$ is a hyperparameter that controls the noise strength. Diffusion model then learns a function $f_{\theta}(\bm{x}_{t},t)$, typically modeled by a U-Net [25], to estimate the $\bm{\epsilon}_{t}$ by minimizing the diffusion loss:
$\mathcal{L}_{diffusion}=\mathbb{E}_{\bm{x}_{0},\bm{\epsilon}_{t},t}\parallel\bm{\epsilon}_{t}-f_{\theta}(\bm{x}_{t},t)\parallel_{2}^{2}.$ (2)
After training, a new data $\bm{x}_{0}$ can be generated from the random noise $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$ through the iterative sampling process, which is formulated as:
$\bm{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha_{t}}}}\bm{\epsilon}_{t})+\sigma_{t}\bm{z},$ (3)
where $\bm{\epsilon}_{t}$ is estimated by the learned function $f_{\theta}(\bm{x}_{t},t)$, $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$, and $\sigma_{t}$ is usually set to $\beta_{t}$ or scaled form of $\beta_{t}$.
More recently, to have a control over the denoising process and generate data of interest, conditional diffusion model with classifier-free guidance is often used because of its efficiency [5]. In particular, a condition $y$ is fed into $f_{\theta}$ with a certain probability during training to get the conditional estimation of the noise $\bm{\epsilon}_{t}$. During sampling, the noise $\bm{\epsilon}_{t}$ is estimated by $(1+w)f_{\theta}(\bm{x}_{t},t,y)-wf_{\theta}(\bm{x}_{t},t,y=\phi)$ with the weight $w$ being set to balance the conditional and unconditional estimations.
@@ 3.2 BEVDiffuser with Ground-Truth Guidance
Training. In the absence of ground-truth BEV feature map $\bm{x}_{0}$, we add noise $\bm{\hat{\epsilon}}_{t}$ to the predicted BEV $\bm{x}_{t_{0}}(0\leq t_{0}\ll T)$ to get $\bm{x}_{t}$. In this case, we don’t have access to the true noise $\bm{\epsilon}_{t}$, which is supposed to be added to $\bm{x}_{0}$ to generate $\bm{x}_{t}$. As a result, instead of using $f_{\theta}$ to estimate the unknown $\bm{\epsilon}_{t}$, we propose to optimize $f_{\theta}$ towards $\bm{x}_{0}$. Since $\bm{x}_{t_{0}}$ is already a good estimation of $\bm{x}_{0}$ with bounded task errors, we first optimize $f_{\theta}$ towards $\bm{x}_{t_{0}}$ by minimizing the diffusion loss $\mathcal{L}_{diffusion}$ defined in Equation 4. To further improve the estimation accuracy, we attach task heads to consume the outputs of $f_{\theta}$ and generate task-specific predictions. In this way, $f_{\theta}$ can also be optimized through the task-specific loss $\mathcal{L}_{task}$. To sum up, we adopt the weighted sum of both losses as the overall loss $\mathcal{L}_{total}^{\scriptscriptstyle diff}$ to train $f_{\theta}$. Equation 5 defines the loss $\mathcal{L}_{total}^{\scriptscriptstyle diff}$ where $\lambda$ denotes a weight.
@@
Sampling. We adopt classifier-free guidance in sampling process where we interpolate between conditional and unconditional outputs of $f_{\theta}$ to get the final estimation of $\bm{x}_{0}$ as Equation 6 calculates. The unconditional estimation of $\bm{x}_{0}$ is obtained by replacing the conditioning layout $l$ with the empty layout $l_{\phi}=\{o_{0},o_{p},...,o_{p}\}$ that only contains points $o_{p}$. We then derive $\bm{\epsilon}_{t}$ from the estimated $\bm{x}_{0}$ by Equation 7 for the iterative sampling process (Equation 3).
@@ 4 Experiments
We validate BEVDiffuser on 3D object detection task, the most common task used to evaluate the effectiveness of the learned BEV feature maps [15, 19, 12, 41]. 3D object detection is critical in autonomous driving that requires both semantic and geometric understanding of the environment to identify and locate objects in 3D space. In this section, we first introduce our experimental setting in Sec. 4.1. In Sec. 4.2, we showcase the capacity of BEVDiffuser in denoising and generating BEV feature maps. We further demonstrate plug-and-play performance of BEVDiffuser in Sec. 4.3 by comparing BEVDiffuser enhanced BEV models with their baseline counterparts.
@@ 4.3 Plug-and-Play Performance of BEVDiffuser
It is worth highlighting that BEVDiffuser brings performance enhancement to BEV models at no cost of any additional adaptation efforts or computational overhead. As a training-only plug-in, BEVDiffuser is removed at deployment, leaving an enhanced BEV model with the architecture unchanged, which is then used for testing. As a result, our BEVDiffuser enhanced models share the same network size and latency as their baseline counterparts which are summarized in Tab. 2. Unlike previous work [12, 41] that need to pass their integrated diffusion models multiple times to denoise the BEV feature maps on-the-fly, our method is more flexible and superior in latency-critical applications like autonomous driving.
@@ 5 Conclusion and Future Work
In this work, we present BEVDiffuser, a novel diffusion model that denoises BEV feature maps using ground-truth guidance. BEVDiffuser consists of a U-Net model trained on BEV feature maps generated by existing BEV models. The U-Net model predicts clean BEV feature maps conditioned on the ground-truth object layout, which then derives the denoising process. BEVDiffuser can be used as a training-only plug-and-play module to enhance the existing BEV models by providing denoised BEV feature maps as additional supervision to BEV predictions. Extensive experiments on challenging nuScenes dataset demonstrate BEVDiffuser’s exceptional denoising and generation capabilities, resulting in significant improvements to existing BEV models, without the need for architectural changes or additional computational overhead. Moreover, results on long-tail object detection and under challenging weather and lighting conditions further confirm the efficacy of BEVDiffuser in improving the BEV quality. In future work, we plan to investigate potential applications of BEVDiffuser for other autonomous driving tasks, such as motion prediction and data augmentation for corner cases.
@@ 7 Implementation Details
8 Ablation Study
We conduct an ablation study on BEVDiffuser ($\mathrm{BD}^{tiny}$) to validate our design choices of layout conditioning and optimization objective, i.e. optimizing towards $x_{t_{0}}$ with the task loss. Note that to optimize towards $\hat{\epsilon}_{t}$, we are not able to attach the task head or use the task loss. As shown in Tab. 5, without the task loss, whether we optimize towards $x_{t_{0}}$ or $\hat{\epsilon}_{t}$, the denoising capability we obtained is quite limited, demonstrating that the task loss is critical to guarantee the denoising performance. Similarly, our layout conditioning also contributes to the superior denoising capability of BEVDiffuser, as evidenced by the inferior performance of the unconditional model.
# denoising steps
Method obj. 1 3 5 10
Ours $x_{t_{0}}$ 35.8/47.7 40.4/52.3 40.8/52.7 40.3/52.3
--task $x_{t_{0}}$ 24.5/34.7 23.1/32.8 21.7/31.0 17.4/26.1
$\hat{\epsilon}_{t}$ 25.2/35.5 25.2/35.5 25.2/35.5 25.2/35.5
--cond. $x_{t_{0}}$ 25.4/35.4 25.3/35.3 25.1/35.0 24.7/34.6
Table 5: Ablation study. mAP/NDS achieved by the variants of BEVDiffuser ($\mathrm{BD}^{tiny}$) with increasing denoising steps (1$\rightarrow$10). Results validate that both the task loss and the layout conditioning contribute to the superior denoising capability of BEVDiffuser.